OpenStudy (anonymous):

If X and Y are independent random variables following N(8,2) and N(12,4*root3) respectively. Find value of 'a' such that P (2X-Y<=a^2) = P (X + 2Y>= a)

OpenStudy (amistre64):

normally distributed with (mean.variance) eh

OpenStudy (anonymous):

oh but whats the method to solve?

OpenStudy (amistre64):

i cant really say that im familiar with the 2x-y setup to determine. would we go with a 95% spread or just means or ... i cant make sense of that part

OpenStudy (anonymous):

well I dont know anything about.. :( just tell me whatever you think the way to solve this

OpenStudy (amistre64):

if i were to take a guess, id think (and i could be wrong on this), that the 3sd spread is what the are relating this to ....

OpenStudy (amistre64):

X = (8+3sqrt(2)) - (8-3sqrt(2)) = 6(2)^(1/2) Y = (12+3(sqrt(4rt3))) - (12-3sqrt(4rt3)) = 12(3)^(1/4)

OpenStudy (amistre64):

then its just mathing out the innards and comparing results, but thats just a guess is all

OpenStudy (amistre64):

i see some content in the internet that suggest to subtract means and add variances .... but im unclear as to why

OpenStudy (anonymous):

o.O but I didnt get it, whatever you have done. Basically it is related with the area under normal distribution . and that 2X+Y and X+2Y are the lines. The area under normal curve before or after these lines as per the sign must be found out and equated. and for that we have to find Standard normal variate Z. I dont know how to find it out in case of 2 variables X and Y. but for single variable it is given byt Z= (X - m) / SD where m=mean and SD= standard deviation

OpenStudy (amistre64):

are there, or was there, any options to choose from?

OpenStudy (amistre64):

ive got a site here that says: the mean of the sum of the distribution of random variables is the sum of the means, u(x+y) = u(x)+u(y) the variance is the difference var(x+y) = var(x)-var(y) in this manner we can setup a distribution: Z = X +- Y, N(u(x+-y),var(x-+y))

OpenStudy (amistre64):

2X = X+X, so maybe Z1 = 2X - Y = ((8+8)-12,(2-2)+4rt(3)) = (4,4rt(3)) Z2 = X + 2Y = (8+(12+12),2-(4rt(3)-4rt(3))) = (32,2)

OpenStudy (anonymous):

yep that should be correct

OpenStudy (anonymous):

I found one similar problem in my book. in that for standard deviation they took sqrt od multi squares and adding those for 2 ie. sd of Z1=root (4*4 + 1* 16*3)=root64 =8

OpenStudy (anonymous):

sd for Z2 = root(1*4 + 4*16*3) =root 196=14

OpenStudy (amistre64):

i wish i could say this makes sense now, but im still on the fence ....

OpenStudy (anonymous):

still I cant find solution for a which is required.

OpenStudy (amistre64):

whats the title of your textbook?

OpenStudy (anonymous):

its applied maths 4 by G. V. Kumbhojkar

OpenStudy (amistre64):

let me do some reading, if noone else can determine an answer ill see what i can learn on it

OpenStudy (anonymous):

okay thats fine. actually I think we have reached upto mean and variance correctly. only we have to find standard normal variate by snv= (U-m)/ sd. In the textbook problem which i was looking at they have taken the result as U I mean in this case a^2 and a and found the area in between. but here we cant find snv as a is unknown and then cant find area as snv is unknown. Pretty confusing. hope you'll get the answer.

OpenStudy (amistre64):

https://onlinecourses.science.psu.edu/stat414/node/172 this looks very promising to me

OpenStudy (amistre64):

X = N(8,2) Y = N(12,4*root3) 2X = (2(8),2^2(2)) -Y = (-1(12),(-1)^2(4 rt(3))) ------------------------- = ( 4 , 8 + 4 rt(3) ) for the distribution setup \[Z_1\le\frac{a^2-4}{\sqrt{8+4\sqrt3}}\]

OpenStudy (amistre64):

X + 2Y>= a X = N(8,2) Y = N(12,4*root3) X = (1(8),1^2(2)) 2Y = (2(12),(2)^2(4 rt(3))) ------------------------- = ( 32 , 2+16 rt(3) ) \[Z_2\ge\frac{a-32}{\sqrt{2+16\sqrt3}}\]

OpenStudy (amistre64):

since these are spose to equate, then Z1 = Z2, when a=?

OpenStudy (amistre64):

\[a^2\sqrt{2+16\sqrt{3}}-4\sqrt{2+16\sqrt{3}} = a\sqrt{8+4\sqrt{3}}-32\sqrt{8+4\sqrt{3}}\] lets clean this up alittle \[a^2k-4k = am-32m\] \[a^2k-am+32m-4k = 0\] and using the quadratic formula to solve for a \[a=\frac{m\pm\sqrt{m^2-4(k)(32m-4k)}}{2k}\]

OpenStudy (amistre64):

make sure the sds are correct, and then that how i would solve for "a"

OpenStudy (anonymous):

@amistre64 everything is correct upto mean but for variance we have to take variance= ∑Ci^2 *σi ^2 which you havent taken. then sd= sqrt(variance). and finally we are not suppose to equate Z1 and Z2 but their probabilities which are depending on the value of a.

OpenStudy (anonymous):

for variance you have taken only Ci^2 and not σi^2.

OpenStudy (amistre64):

o^2 is variance, and variance is given in the notation:\[N(\mu,\sigma^2)\] simga^2 is variance

OpenStudy (amistre64):

library is closing early today, so im outta time :/

OpenStudy (anonymous):

oh thats fine not a problem. Thanks a lot for your support we'll solve this later.

OpenStudy (amistre64):

Tapping on my phone. But after thinking last night we are creating 2 new normal disteibutions W and T we then take and standardize them both to Z in order to compare their probabilities. The symmetry of a normal distribution is such that the left and right tails of -z and z are equal.